# Corvus A: A low-mass, isolated galaxy at 3.5 Mpc

### v1

@@ 2.1 Magellan Megacam imaging
We perform point-spread function (PSF) fitting photometry on the final image stacks, using the DAOPHOT and ALLFRAME software suite (Stetson, 1987, 1994), following the same methodology described in Mutlu-Pakdil et al. (2021). We remove objects that are not point sources by culling our catalogs of outliers in versus magnitude, magnitude error versus magnitude, and sharpness versus magnitude space. Instrumental magnitudes are then calibrated to the DECaLS DR10 catalog. We correct for Galactic extinction on a star-by-star basis using the Schlegel et al. (1998) reddening maps with the coefficients from Schlafly & Finkbeiner (2011). The resulting color magnitude diagram (CMD) of Corvus A is shown in Figure 2, which is discussed further in §3.2 & §4.1.
@@ 3.5 Star formation rate and stellar mass estimates
Although Corvus A was detected in all three Swift/UVOT bands, we use only UVM2 to determine the star formation rate (SFR) as this is most similar to GALEX NUV (Hoversten et al., 2009). The UV flux was measured within an elliptical aperture with a semi-major axis equal to two half-light radii (Table 1). This resulted in a UV magnitude of . This was then converted to a SFR following the relations of Iglesias-Páramo et al. (2006) for NUV, giving .
To estimate the total stellar mass of Corvus A we use its color and magnitudes in the mass-to-light ratio relations of Zibetti et al. (2009) and Into & Portinari (2013). These give stellar mass estimates of and 6.04, respectively. We take the log mean of these values as Corvus A’s stellar mass and their difference as an approximate indication of the uncertainty, i.e. .222Note that the distance uncertainty would contribute an additional 0.06 dex to the total uncertainty in the mass. We have not incorporated distance uncertainties in the uncertainty estimates of any absolute quantities in Table 1. As with its H i mass this is a few times more massive than Leo P (or Pavo).
As the young stars are dominating the light from Corvus A, it is quite likely that these mass-to-light ratio relations have under-counted the contribution from Corvus A’s old stellar population. A more accurate estimate of the total stellar mass will be possible with the recent HST observations of Corvus A, which fully resolve its stellar population (Mutlu-Pakdil et al. in prep.).

### v2

@@ 2.1 Magellan Megacam imaging
We perform point-spread function (PSF) fitting photometry on the final image stacks, using the DAOPHOT and ALLFRAME software suite (Stetson, 1987, 1994), following the same methodology described in Mutlu-Pakdil et al. (2021). We remove objects that are not point sources by culling our catalogs of outliers in versus magnitude, magnitude error versus magnitude, and sharpness versus magnitude space. Instrumental magnitudes are then calibrated to the DECaLS DR10 catalog. We correct for Galactic extinction on a star-by-star basis using the Schlegel et al. (1998) reddening maps with the coefficients from Schlafly & Finkbeiner (2011), which were evaluated according to the Fitzpatrick (1999) reddening law with normalization . Specifically, we adopted the coefficients of 3.237 and 2.176 for and to convert E(B-V) to and , respectively. The resulting color magnitude diagram (CMD) of Corvus A is shown in Figure 2, which is discussed further in §3.2 & §4.1.
@@ 3.5 Star formation rate and stellar mass estimates
Although Corvus A was detected in all three Swift/UVOT bands, we use only UVM2 to determine the star formation rate (SFR) as this is most similar to GALEX NUV (Hoversten et al., 2009). The UV flux was measured within an elliptical aperture with a semi-major axis equal to two half-light radii (Table 1). This resulted in a UV magnitude of . This was then converted to a SFR following the relations of Iglesias-Páramo et al. (2006) for NUV, giving (including the distance uncertainty).
To estimate the total stellar mass of Corvus A we use its color and magnitudes in the mass-to-light ratio relations of Zibetti et al. (2009) and Into & Portinari (2013). These give stellar mass estimates of and 6.04, respectively. We take the log mean of these values as Corvus A’s stellar mass and their difference as an approximate indication of the uncertainty, i.e. . As with its H i mass this is a few times more massive than Leo P (or Pavo). As the young stars are dominating the light from Corvus A, it is quite likely that these mass-to-light ratio relations have under-counted the contribution from Corvus A’s old stellar population and that a more in-depth characterization of its stellar population would result in a slightly higher stellar mass estimate.